# Generalized saddle-node ghosts and their composite structures in dynamical systems

### 09 Apr 2026 v2

@@ 1 Introduction
In this article, we address this gap for one of the possible mechanisms underlying long transients and sequential transitions between them: ghost states and their composite structures. By giving a novel generalized definition of ghost states, we provide a new theoretical framework that allows to better describe ghosts as phase space objects and phenomena associated with ghosts, many of which have not been reported before. Based on this framework, we introduce a series of novel algorithms for identifying and characterizing ghost states and their composite structures (e.g., ghost channels) along system trajectories, enabling us to obtain new insights into the transient dynamics, e.g., of neuronal and gene regulatory network models and to discover novel transient dynamics phenomena such as bifurcations between ghosts.
@@ 2.1 Theoretical framework
Visualizing the dynamics by plotting $\dot{x}$ vs $x$, we immediately see that the system has two fixed points, one at $x_{s}^{*}=-\sqrt{\mu}$ and one at $x_{u}^{*}=\sqrt{\mu}$ (Figure 1a). Inferring the direction of flow by considering the value of $\dot{x}$ in the vicinity of $x_{s}^{*}$ (black arrowheads in Figure 1a), we can see that trajectories will move towards $x_{s}^{*}$, hence $x_{s}^{*}$ is a stable fixed point. By the same kind of reasoning, we find that $x_{u}^{*}$ is unstable. In higher dimensions, sinks and repellers exist, too, but are complemented by saddles, defined to have one repelling and one attracting direction on the plane, or $k$ repelling and nkn-k attracting directions in $n$ dimensions, respectively (Figure 1b).
@@
For $j=1$, Definition 2.1 is equivalent to the definitions given in [3, 4], but for $j>1$ captures the other saddle-nodes shown in Figure 1d: (SN1) ensures a center subspace space spanned by $j$ independent eigendirections, (SN2) means that for each center direction, there is a parameter that upon perturbation leads to an unfolding of the saddle-node, i.e. saddle-nodes of type $j,k$ are structurally unstable. Note that since this can include different parameters for different center directions, the saddle-node of type $j,k$ has a codimension of $\leq j$. (SN3) guarantees quadratic non-degeneracy for each non-zero center direction. (SN4) simply describes the dimension of the saddle-nodes’ stable and unstable subspaces.
Having defined saddle-nodes of type $j,k$, we next will consider how their ghost influences the non-asymptotic dynamics of a system after a SN-bifurcation occurs. To do so, we return again to Equation 1, this time for a small $\mu>0$ (Figure 1e). Since $\dot{x}=\mu+x^{2}>0$ for $\mu>0$, Equation 1 has no fixed points anymore, i.e., no invariant objects can govern the system’s dynamics. However, while $\dot{x}>$, $\mu$ is small and so for small $|x|$, $\dot{x}$ becomes small, too, and any trajectory starting on the negative half-line spends a long time in the vicinity of 0 before eventually escaping towards infinity (cf. inset in Figure 1e). Notice how the phase portrait of the ghost, except for the vanished fixed point, does not differ much from that of it’s saddle-node at $\mu$, in particular the direction of the flow remains the same. Indeed, this simple example already shows many features that are associated with ghosts: slow dynamics, non-invariance and a finite-time sense of attraction. However, as saddle-nodes can become more complex in higher dimensions, so do their ghosts. Consider the following system:
@@
While we have not explicitly stated how the set $\mathcal{A}$ is to be defined, we find in practice a small neighborhood of an expected ghost to be sufficient. (G1) formally captures the slow dynamics that result from the flat quasi-potential landscape due to the coalescence of equilibria at the saddle-node [62, 38, 37]. The auxiliary function $Q$ is equivalent to the cost-function $J$ used in [77]. (G2) simply states the non-invariance of ghosts, i.e., any trajectory visiting the neighborhood of a ghost must eventually leave this neighborhood again. While (G1-G2) are necessary conditions for ghosts, they are likely met by most long transients. (G3) is the core of the definition and states that each ghost has a closest saddle-node of type $j,k$ in parameter space from which it inherits the organization of the phase portrait and which defines the ghost’s dimension. It also excludes that a ghost can be a ghost of multiple saddle-nodes and will be important in the context of different ways of unfolding a type $j,k$ saddle-node. Although the notion of attraction is only implicit (as it is not necessary for the purpose of this study), definition 2.2 accounts for all features observed for the ghosts highlighted above.
@@ 2.2 The GhostID algorithm
Having formally defined what we mean by a ghost brings us a step closer to an algorithm that can identify and characterize ghosts. While (G1) and (G2) are straightforward to test, searching the parameter space for parental saddle-nodes is not a feasible strategy. We thus need move on from (G3) to an algorithmically identifiable criterion. For this, we will make use of our previous observation [38] that all ghosts we examined so far exhibited a linear spatial distribution of instantaneous eigenvalues ranging from negative to positive along the former center direction of the saddle-node. Indeed, we find the same to be true for higher dimensional ghosts (e.g., for system 1.7 for $n=j=2,k=0$, cf. Supplementary Figure 1). This leads us to formulate the following conjecture:
@@
Conjecture 2.1 states that the instantaneous eigenvalue distributions along the former center directions of the saddle-node still persist at the ghost, which is what we will use as algorithmic criterion to evaluate if a trajectory passes nearby a ghost. Although we were not able to provide a proof here, we strongly suspect conjecture 2.1 follows from definition 2.1 and 2.2 and find the existence of such a distribution in all systems exhibiting a ghost state resulting from a SN-bifurcation. While such a spatial distribution of instantaneous eigenvalues might potentially result from other mechanisms, too, and hence does not strictly guarantee the presence of a ghost, we empirically find it to be a very strong indicator based on numerous tested systems with different mechanisms underlying long transients.
@@
Following the outlined framework, we propose a simple algorithm, which we term GhostID, that identifies ghosts along a trajectory by evaluating criteria (G1), (G2) and the changes in instantaneous eigenvalues from conjecture 2.1. The basic steps of this algorithm, illustrated in Figure 2, are as follows: The first step of the algorithm is to find the locally slowest points along a trajectory, corresponding to evaluating a trajectory proxy for criterion (G1). In the second step, the algorithm evaluates if the trajectory leaves the $\varepsilon$-neighborhood of a locally slowest point, testing for non-invariance (G2). The third and final step consists of testing whether one or more of the instantaneous eigenvalues change from negative to positive along the trajectory segment within the $\varepsilon$-neighborhood from the second step. Thus, starting from a slow point, if both non-invariance and the eigenvalue criterion are satisfied, we consider the long transient to come from a ghost, with the number of eigenvalues changing sign being considered by us as the ghost’s dimension. A full technical description of the algorithm, its hyper-parameters and some rationales for choosing their values them can be found in Supplementary section 1.1 and Supplementary Figure 2a-b.
@@ 2.3 Numerical validation of the GhostID algorithm
where $C$, $M$ and $B$ represent corals, macroalgae and benthic r-strategists (groups of algae that colonize disturbed reefs on a faster timescale than macroalgae), and parameters $a,b,r$ and $\gamma$ represent (over-)growth rates, $g$ the grazing rate by herbivores, and $m$ is the mortality rate of corals. This model has been proposed to be organized close to a saddle-node bifurcation where it exhibits a ghost attractor to explain how coral reef recovery may be hampered by repeated climate or human driven perturbations that prevent trajectories escaping a long transient of the algae-dominated state [8]. Applying GhostID to a trajectory from this regime correctly identifies a ghost between the system’s nullclines (Figure 3a, left). Considering the pQpQ time series reveals that the algorithm interestingly identified two $Q$-minima along the trajectory (Figure 3a, middle). Once the trajectory reaches the system’s stable fixed point, $pQ(t)$ appears to fluctuate irregularly, leading to additional $Q$-minima. However, these irregular movements are mere numerical noise which can be reduced (but not eliminated due to limited floating point precision) by lowering the numerical tolerances of the integration procedure (Supplementary Figure 4) and have been ignored by GhostID by requiring peaks identified by SciPy’s find_\_peaks to have a minimum width of 5050dt. The eigenvalues along the trajectory segment around the first identified $Q$-minimum are consistent with the system visiting a slow region around the saddle on the y-axis and are thus not identified as a ghost, whereas on the segment of the second minimum eigenvalue $\lambda_{2}$ changes sign from negative to positive, leading to the algorithm identifying a ghost (Figure 3a, right). Additional examples of GhostID correctly identifying known ghosts in models from cellular biochemistry, gene regulatory networks, and condensed matter physics can be found in Supplementary section 1.4.1 and Supplementary Figure 5. Lastly, while we are not aware of higher-dimensional or non-attracting ghosts being described in the literature, GhostID correctly identifies these ghosts and their corresponding dimensions from our example system 1.7 (Supplementary Figure 6).
@@
where NN denotes the prey species, $P$ the predators, and $\alpha$ is the growth rate of the prey population, $K$ its carrying capacity, $\gamma$ the predation rate, $h$ the half-saturation constant for predation, $v$ is a scaling factor of how predation influences predator reproduction, $m$ is the mortality rate of predators and $\epsilon$ the timescale separation between predator and prey dynamics. For the parameter regime depicted in Figure 3b (left), the system exhibits oscillatory dynamics in which the orbit is forced to pass close to two different saddles (saddle crawl-by), causing the trajectory to slow down in their vicinity. However, the eigenvalues along the trajectory segments of the corresponding $Q$-minima do not cross 0, hence the saddle crawl-bys are not identified as ghosts by our algorithm (Figure 3b, middle and right). Another example involving a heteroclinic cycle can be found in Supplementary section 1.4.2 and Supplementary Figure 7. Since saddles are by definition hyperbolic fixed points, it is easy to prove that within a sufficiently small environment of the fixed point, the eigenvalues along a trajectory in this environment can never change sign (Supplementary section 1.5) and thus will not be identified as ghosts by our algorithm.
@@ 2.4.1 Two-dimensional ghosts and ghost-bifurcations in coupled theta neurons
To begin our investigation into complex and potentially novel ghost dynamics, we decided to focus on simple models of coupled neurons whose dynamics in isolation exhibit a SNIC bifurcation as key ingredient for the emergence of ghosts. While long transients and metastability in computational neuroscience have received some attention over time (cf. [25, 54] for reviews), the role of ghosts in neuronal dynamics remains relatively underexplored. A good example for the importance of ghosts is a recent study by Choi et al. which proposed fly olfactory neurons to be organized close to a SNIC bifurcation, showing that the resulting ghost dynamics provide optimal information processing for sensory stimuli [13]. We thus consider the theta neuron [19] as a simple phase oscillator model that captures the essential dynamic features described in [13] to explore how pulse coupling of two identical theta neurons influences the ghost dynamics. The equations for this systems are given by [5]:
@@
Figure 4: Higher-dimensional ghosts in coupled theta neuron models. a, left: phase space organization of identical theta neurons before bifurcation point. Black dot indicates sink, white dot indicates repeller, grey dots indicate saddles. Right: trajectory and two-dimensional ghost identified by GhostID after bifurcation. a,d White lines indicate trajectories, grey arrows flow. Note that for better visualization, $\theta_{1}$ and $\theta_{2}$ have been phase-shifted by $\pi$ in these plots, such that the origin is shown at $(\pi,\pi)$. b, tracking of the ghost identified in a in a bifurcation diagram using the track_\_ghost_\_branch function. Each ghost along the branch is color-coded by its trapping duration. c, two-parameter bifurcation diagram for non-identical Theta-neuons with distinct excitability thresholds. Boundary between areas with one- and two-dimensional ghosts indicate ghost SN-bifurcation (gSN). Red line indicates SNIC bifurcation. d, phase portraits for parameters yielding one-dimensional ghosts.
Next, we turn to the case of non-identical neurons, i.e., allowing for $\eta_{1}\neq\eta_{2}$. While the track_\_ghost_\_branch currently does feature tracking of ghosts versus two parameters, we can make use of the ghostID_\_phaseSpaceSample method to characterize the ghost dynamics across the $\eta_{1}$-$\eta_{2}$-plane from a Latin hypercube sample within a specified region in phase space. Using this approach, we find two transitions from a parameter region in which only fixed points exist (i.e. without any ghosts) towards regions in which only one-dimensional ghosts exist after crossing a SNIC bifurcation (Figure 4c). Both SNIC branches merge at $\eta_{1}=\eta_{2}=0$, corresponding to two saddle-nodes of type $1,0$ coalescing at $\theta_{1}=\theta_{2}=0$ to form a type $2,0$ saddle-node. After passing one of the SNIC bifurcations, the system exhibits a stable and an unstable limit cycle due to one of the neurons still being in an excitable, not spiking regime (Figure 4d). Interestingly, moving from the 1-dimensional ghost region into the region defined by both $\eta_{1},\eta_{2}>0$ reveals another transition into a region where two-dimensional ghosts exist. This transition seems to be best described as a ghost SN-bifurcation, in which the attracting one-dimensional ghost merges with the non-attracting one-dimensional ghost shown in (Figure 4d). Indeed, we observe the same bifurcation between the ghost of a type $1,0$ saddle-node and the ghost of a type $1,1$ saddle-node for system 1.7 (Supplementary Figure 9). Together these results highlight the capability of GhostID to detect ghosts, track them across parameter space and identify transitions in their properties.

### 16 Apr 2026 v3

@@ 1 Introduction
In this article, we address this gap for one of the possible mechanisms underlying long transients and sequential transitions between them: ghost states and their composite structures. By giving a novel generalized definition of ghost states, we provide a new theoretical framework that allows to better describe ghosts as phase space objects and phenomena associated with ghosts, many of which have not been reported before. Based on this framework, we introduce a series of novel algorithms for identifying and characterizing ghost states and their composite structures (e.g., ghost channels [38]) along system trajectories, enabling us to obtain new insights into the transient dynamics, e.g., of neuronal and gene regulatory network models and to discover novel transient dynamics phenomena such as bifurcations between ghosts.
@@ 2.1 Theoretical framework
Visualizing the dynamics by plotting $\dot{x}$ vs $x$, we immediately see that the system has two fixed points, one at $x_{s}^{*}=-\sqrt{\mu}$ and one at $x_{u}^{*}=\sqrt{\mu}$ (Figure 1a). Inferring the direction of flow by considering the value of $\dot{x}$ in the vicinity of $x_{s}^{*}$ (black arrowheads in Figure 1a), we can see that trajectories will move towards $x_{s}^{*}$, hence $x_{s}^{*}$ is a stable fixed point. By the same kind of reasoning, we find that $x_{u}^{*}$ is unstable. In higher dimensions, sinks and repellers exist, too, but are complemented by saddles, defined to have one repelling and one attracting direction on the plane, or $k$, $0<k<n$, repelling and nkn-k attracting directions in $n$ dimensions, respectively (Figure 1b).
@@
For $j=1$, Definition 2.1 reduces to the definitions given in [3, 4], but for $j>1$ captures the other saddle-nodes shown in Figure 1d: (SN1) ensures a center subspace space spanned by $j$ independent eigendirections, (SN2) means that for each center direction, there is a parameter that upon perturbation leads to an unfolding of the saddle-node, i.e. saddle-nodes of type $j,k$ are structurally unstable. Note that since this can include different parameters for different center directions, the saddle-node of type $j,k$ has a codimension of $\leq j$. (SN3) guarantees quadratic non-degeneracy for each non-zero center direction. (SN4) simply describes the dimension of the saddle-nodes’ stable and unstable subspaces.
Having defined saddle-nodes of type $j,k$, we next will consider how their ghost influences the non-asymptotic dynamics of a system after a SN-bifurcation occurs. To do so, we return again to Equation 1, this time for a small $\mu>0$ (Figure 1e). Since $\dot{x}=\mu+x^{2}>0$ for $\mu>0$, Equation 1 has no fixed points anymore, i.e., no invariant objects can govern the system’s dynamics. However, while $\dot{x}>0$, $\mu$ is small and so for small $|x|$, $\dot{x}$ becomes small, too, and any trajectory starting on the negative half-line spends a long time in the vicinity of 0 before eventually escaping towards infinity (cf. inset in Figure 1e). Notice how the phase portrait of the ghost, except for the vanished fixed point, does not differ much from that of its saddle-node at $\mu$, in particular the direction of the flow remains the same. Indeed, this simple example already shows many features that are associated with ghosts: slow dynamics, non-invariance and a finite-time sense of attraction. However, as saddle-nodes can become more complex in higher dimensions, so do their ghosts. Consider the following system:
@@
While we have not explicitly stated how the set $\mathcal{A}$ is to be defined, we find in practice a small neighborhood of a potential ghost to be sufficient. (G1) formally captures the slow dynamics that result from the flat quasi-potential landscape due to the coalescence of equilibria at the saddle-node [62, 38, 37]. The auxiliary function $Q$ is equivalent to the cost-function $J$ used in [77]. (G2) simply states the non-invariance of ghosts, i.e., any trajectory visiting the neighborhood of a ghost must eventually leave this neighborhood again. While (G1-G2) are necessary conditions for ghosts, they are likely met by most long transients. (G3) is the core of the definition and states that each ghost has a closest saddle-node of type $j,k$ in parameter space from which it inherits the organization of the phase portrait and which defines the ghost’s dimension. It also excludes that a ghost can be a ghost of multiple saddle-nodes and will be important in the context of different ways of unfolding a type $j,k$ saddle-node. Although the notion of attraction is only implicit (as it is not necessary for the purpose of this study), definition 2.2 accounts for all features observed for the ghosts highlighted above.
@@ 2.2 The GhostID algorithm
Having formally defined what we mean by a ghost brings us a step closer to an algorithm that can identify and characterize ghosts. While (G1) and (G2) are straightforward to test, searching the parameter space for parental saddle-nodes is not a feasible strategy. We thus need move on from (G3) to an algorithmically identifiable criterion. For this, we will rely on our previous observation [38] that all ghosts we examined so far exhibit a linear spatial distribution of instantaneous eigenvalues ranging from negative to positive along the former center direction of the saddle-node. Indeed, we find the same to be true for higher dimensional ghosts (e.g., for system 1.7 for $n=j=2,k=0$, cf. Supplementary Figure 1). This leads us to formulate the following conjecture:
@@
Conjecture 2.1 states that the instantaneous eigenvalue distributions along the former center directions of the saddle-node still persist at the ghost, which is what we will use as algorithmic criterion to evaluate if a trajectory passes nearby a ghost. Although we were not able to provide a proof here, we strongly suspect conjecture 2.1 follows from definition 2.1 and 2.2 and find empirically that such a distribution exists in all systems exhibiting a ghost state resulting from a SN-bifurcation. While such a spatial distribution of instantaneous eigenvalues might potentially result from other mechanisms, too, and hence does not strictly guarantee the presence of a ghost, we empirically find it to be a very strong indicator based on numerous tested systems with different mechanisms underlying long transients.
@@
Following the outlined framework, we propose a simple algorithm, which we term GhostID, that identifies ghosts along a trajectory by evaluating criteria (G1), (G2) and the changes in instantaneous eigenvalues from conjecture 2.1. The basic steps of this algorithm, illustrated in Figure 2, are as follows: The first step of the algorithm is to find the locally slowest points along a trajectory, corresponding to evaluating a trajectory proxy for criterion (G1). In the second step, the algorithm evaluates if the trajectory leaves the $\varepsilon$-neighborhood of a locally slowest point, testing for non-invariance (G2). The third and final step consists of testing whether one or more of the instantaneous eigenvalues change from negative to positive along the trajectory segment within the $\varepsilon$-neighborhood from the second step. Thus, starting from a slow point, if both non-invariance and the eigenvalue criterion are satisfied, we consider the long transient to be caused by a ghost, with the number of eigenvalues changing sign being considered by us as the ghost’s dimension. A full technical description of the algorithm, its hyper-parameters and some rationales for choosing their values them can be found in Supplementary section 1.1 and Supplementary Figure 2a-b.
@@ 2.3 Numerical validation of the GhostID algorithm
where $C$, $M$ and $B$ represent corals, macroalgae and benthic r-strategists (groups of algae that colonize disturbed reefs on a faster timescale than macroalgae), and parameters $a,b,r$ and $\gamma$ represent (over-)growth rates, $g$ the grazing rate by herbivores, and $m$ is the mortality rate of corals. This model has been proposed to be organized close to a saddle-node bifurcation where it exhibits a ghost attractor to explain how coral reef recovery may be hampered by repeated climate or human driven perturbations that prevent trajectories escaping a long transient of the algae-dominated state [8]. Applying GhostID to a trajectory from this regime correctly identifies a ghost between the system’s nullclines (Figure 3a, left). Considering the $pQ$ time series reveals that the algorithm interestingly identified two $Q$-minima along the trajectory (Figure 3a, middle). Once the trajectory reaches the system’s stable fixed point, $pQ(t)$ appears to fluctuate irregularly, leading to additional $Q$-minima. However, these irregular movements are mere numerical noise which can be reduced by lowering the numerical tolerances of the integration procedure (Supplementary Figure 4) and have been ignored by GhostID by requiring identified peaks to have a minimum width of 5050dt. The eigenvalues along the trajectory segment around the first identified $Q$-minimum are consistent with the system visiting a slow region around the saddle on the y-axis and are thus not identified as a ghost, whereas on the segment of the second minimum eigenvalue $\lambda_{2}$ changes sign from negative to positive, leading to the algorithm identifying a ghost (Figure 3a, right). Additional examples of GhostID correctly identifying known ghosts in models from cellular biochemistry, gene regulatory networks, and condensed matter physics can be found in Supplementary section 1.4.1 and Supplementary Figure 5. Lastly, while we are not aware of higher-dimensional or non-attracting ghosts being described in the literature, GhostID correctly identifies these ghosts and their corresponding dimensions from our example system 1.7 (Supplementary Figure 6).
@@
where $N$ denotes the prey species, $P$ the predators, and $\alpha$ is the growth rate of the prey population, $K$ its carrying capacity, $\gamma$ the predation rate, $h$ the half-saturation constant for predation, $v$ is a scaling factor of how predation influences predator reproduction, $m$ is the mortality rate of predators and $\epsilon$ the timescale separation between predator and prey dynamics. For the parameter regime depicted in Figure 3b (left), the system exhibits oscillatory dynamics in which the orbit is forced to pass close to two different saddles (saddle crawl-by), causing the trajectory to slow down in their vicinity. However, the eigenvalues along the trajectory segments of the corresponding $Q$-minima do not cross 0, hence the saddle crawl-bys are not identified as ghosts by our algorithm (Figure 3b, middle and right). Another example involving a heteroclinic cycle can be found in Supplementary section 1.4.2 and Supplementary Figure 7. Since saddles are hyperbolic fixed points by definition, it is easy to prove that within a sufficiently small neighborhood of the fixed point, the eigenvalues along a trajectory in this neighborhood can never change sign (Supplementary section 1.5) and thus will not be identified as ghosts by our algorithm.
@@ 2.4.1 Two-dimensional ghosts and ghost-bifurcations in coupled theta neurons
To begin our investigation into complex and potentially novel ghost dynamics, we decided to focus on simple models of coupled neurons whose dynamics in isolation exhibit a SNIC bifurcation as key ingredient for the emergence of ghosts. While long transients and metastability in computational neuroscience have received some attention over time (cf. [25, 54] for reviews), the role of ghosts in neuronal dynamics remains relatively underexplored. A good example for the importance of ghosts was recently provided by Choi et al. which proposed fly olfactory neurons to be organized close to a SNIC bifurcation, showing that the resulting ghost dynamics provide optimal information processing for sensory stimuli [13]. We thus consider the theta neuron [19] as a simple phase oscillator model that captures the essential dynamic features described in [13] to explore how pulse coupling of two identical theta neurons influences the ghost dynamics. The equations for this systems are given by [5]:
@@
Figure 4: Higher-dimensional ghosts in coupled theta neuron models. a, left: phase space organization of identical theta neurons before bifurcation point. Black dot indicates sink, white dot indicates repeller, grey dots indicate saddles. Right: trajectory and two-dimensional ghost identified by GhostID after bifurcation. a,d White lines indicate trajectories, grey arrows flow. Note that for better visualization, $\theta_{1}$ and $\theta_{2}$ have been phase-shifted by $\pi$ in these plots, such that the origin is shown at $(\pi,\pi)$. b, tracking of the ghost identified in a in a bifurcation diagram using the track_\_ghost_\_branch function. Each ghost along the branch is color-coded by its trapping duration. c, two-parameter bifurcation diagram for non-identical theta neruons with distinct excitability thresholds. Boundary between areas with one- and two-dimensional ghosts indicate ghost SN-bifurcation (gSN). Red line indicates SNIC bifurcation. d, phase portraits for parameters yielding one-dimensional ghosts.
Next, we turn to the case of non-identical neurons, i.e., allowing for $\eta_{1}\neq\eta_{2}$. While track_\_ghost_\_branch currently does not feature tracking of ghosts versus two parameters, we can apply the ghostID_\_phaseSpaceSample method to characterize the ghost dynamics across the $\eta_{1}$-$\eta_{2}$-plane from a Latin hypercube sample within a specified region in phase space. Using this approach, we find two transitions from a parameter region in which only fixed points exist (i.e. without any ghosts) towards regions in which only one-dimensional ghosts exist after crossing a SNIC bifurcation (Figure 4c). Both SNIC branches merge at $\eta_{1}=\eta_{2}=0$, corresponding to two saddle-nodes of type $1,0$ coalescing at $\theta_{1}=\theta_{2}=0$ to form a type $2,0$ saddle-node. After passing one of the SNIC bifurcations, the system exhibits a stable and an unstable limit cycle due to one of the neurons still being in an excitable, not spiking regime (Figure 4d). Interestingly, moving from the 1-dimensional ghost region into the region defined by both $\eta_{1},\eta_{2}>0$ reveals another transition into a region where two-dimensional ghosts exist. This transition seems to be best described as a ghost SN-bifurcation, in which the attracting one-dimensional ghost merges with the non-attracting one-dimensional ghost shown in (Figure 4d). Indeed, we observe the same bifurcation between the ghost of a type $1,0$ saddle-node and the ghost of a type $1,1$ saddle-node for system 1.7 (Supplementary Figure 9). Together these results highlight the capability of GhostID to detect ghosts, track them across parameter space and identify transitions in their properties.